OpenStudy (anonymous):

x+2x+5+2x=20 x= 17 right?

OpenStudy (anonymous):

OPPS NVM

OpenStudy (anonymous):

>_________<

OpenStudy (anonymous):

x+2x+5+2x=20 (1+2+2)x + 5 = 20 5x = 20-5 5x = 15 x = ?

OpenStudy (anonymous):

my bad, I WROTE ON MY PAPER 90 INSTEAD OF 20. JUST REALIZED THAT